OpenStudy (anonymous):

Determine the slope and the y-intercept of the graph of the equation. -x + 11y - 33 = 0

OpenStudy (darkprince14):

rewrite the term in ax +by = c

OpenStudy (anonymous):

\[11y=x+33\]

OpenStudy (darkprince14):

you'll get the slope if you transform the equation into the slope form which is y = mx + b..y's coefficient must be 1...

OpenStudy (darkprince14):

the Y coefficient, i mean..

OpenStudy (anonymous):

y intercept=3

OpenStudy (anonymous):

\[y=-\frac{ 1 }{ 11 }x+3\]

OpenStudy (anonymous):

y=1/11x+3 y=-1x+3

OpenStudy (darkprince14):

i think it's positive..

OpenStudy (anonymous):

it can't be \[-1\div11, \]

OpenStudy (darkprince14):

the slope would be.. 1/11...

OpenStudy (darkprince14):

look, -x + 11y = 33 11y = x + 33 if you divide bot sides by 11, you'll get \[y = \frac{ 1 }{ 11 } + 3\]

OpenStudy (darkprince14):

1/11 x

OpenStudy (darkprince14):

forgot to type x

OpenStudy (darkprince14):

\[y = \frac{ 1 }{ 11 }x + 3\]

OpenStudy (darkprince14):

y = mx + b m = slope b = y-intercept

OpenStudy (anonymous):

thanks.